In a constant pressure calorimeter 50.0 mL of 0.310 M Ba(OH)2 was added to 50.0 mL of 0.620 M HCl. the reactions caused the temperature of the solution to rise from 23.77 c to 27.99 c if the solution has the same density and same specific heat as water (1.00 g/mL and 4.184 j/g c ) what is triangle H for this reaction (per mole of H2O produced)

assume that the total volume is the sum of the individual volumes
Answer #1

mass of solution = 50 + 50 = 100 g

moles of Ba(OH)2 = 50.0 x 0.310 / 1000 = 0.0155 mol

moles of HCl = 50.0 x 0.620 / 1000 = 0.031

Q = m Cp dT

   = 100 x 4.184 x (27.99 - 23.77)

   = 1765.65 J

Ba(OH)2   + 2 HCl    ------------> BaCl2   + 2 H2O

    1                  2                              1             2

0.0155          0.031                   

delta H = - Q / n

            = - 1.766 / 0.0155

delta H = - 114 kJ/mol